Part A and B Alternative method The company is developing a greener method of making propene oxide from propene using hydrogen peroxide,
Answer #1

Question 1

As the weight of the final product is given, back calculations need to be made to arrive at the weight of the starting material.

Molar mass of propene oxide = 58.08 g/ mol

Mass of propene oxide = 29.00 kg

Moles of propene oxide = mass/ molar mass = (29.00 x 1000) g / 58.08 g/mol = 499.311 mol

If the yield of the reaction were 100 % then to get 499.311 mol propene oxide, 499.311 mol of propene would be required.

But the yield is 89%. So the amount of propene required is = moles of prod x 100 / yield = 499.311 mol x 100 / 89.0 = 561.024 mol

Molar mass of propene = 42.081 g/ mol

Mass of propene required = moles x molar mass = 561.024 mol x 42.081 g / mol = 23608.45 g = 23.608 kg

Moles of H2O2 = moles of propene = 561.024 mol

(Based on balanced chemical equation)

Molar mass of H2O2 = 34.05 g/mol

Mass of H2O2 = moles x molar mass = 561.024 mol x 34.05 g/mol = 19102.8g = 19.103 kg

Vanadium required 1kg for 1 kg propene. So total vanadium required = 23.608 kg

Vanadium recovered = 75%

So Vanadium consumed = 25%

Mass of vanadium consumed = 25 x 23.608 kg/ 100 = 5.902 kg

H2 O required is 32 Kg per 1 kg of propene

Total H2O required = 32 x 23.608 kg = 755.456 kg

Now the cost of materials used is equal to the quantity of material used in kg x per kg cost of material

Cost of propene = 23.608 kg x $10.97/kg = $258.980

Cost of H2O = 755.456 kg x $1.85/kg = $1397.594

Cost of H2O2 = 19.103 kg x $1.96/kg = $37.442

Cost of vanadium = 5.902 kg x $57.50/kg = $339.365

Total cost for the reaction is sum of all the starting material cost

On addition of above numbers we get cost of $2033.381

Earning on propene oxide = 29 kg x $ 258.25/kg = $7489.25

Profit = $7489.25 - $ 2033.381 = $5455.869

Question 2

Molar mass of propene oxide = 58.08 g/mol

Mass of propene oxide = 29.0 kg

Moles of propene oxide = 29000 g / 58.08 g/mol = 499.311 mol

Yield 96%

So the moles of propene = 499.311 mol x 100 / 96 = 520.116 mol

Mass of propene required = 520.116 mol x 42.081g/mol = 21886.9 g = 21.887kg

Moles of mcpba = 520.116mol

Molar mass of mcpba = 172.56 g/mol

Mass of mcpba = 520.116 mol x 172.56 g/mol = 89751.2 g = 89.751 kg

1 L DCM is required for 21g of propene

For 21887 g propene, DCM required = 1 x 21887/21 = 1042.24 L

Calculating the cost for each material

Cost for propene = 21.887 kg x $10.97 = $240.100

Cost of mcpba = 89.751 kg x $5.28 = $473.885

Cost of DCM = 1042.24 x $ 2.12 = $2209.545

Total cost = $2923.530

Coat of waste = $5.00 x 29 = $145

Final expenses = $2923.530 + $145 = $3068.53

Total earnings = 29 kg x $258.25 = $7489.25

Profit = $ 7489.25 - $3068.530 = $4420.720